OpenStudy (anonymous):

Use a taylor series to approximate definite integral of sin(x^2)dx from x=0 to x=0.2

OpenStudy (anonymous):

fo you know series for sine?

OpenStudy (anonymous):

yes the series is x^2 - x^6/6+x^10/120-x^14/5040+...

OpenStudy (anonymous):

replace \(x\) in that series for \(x^2\) an then take it out as far as you care to series is \[x-\frac{x^3}{3!}+\frac{x^5}{5!}-...\] replace \(x\) by \(x^2\) get \[x^2-\frac{x^6}{3!}+\frac{x^{10}}{5!}-...\]

OpenStudy (anonymous):

just replace \(x\) by \(x^2\) not the denominators

OpenStudy (anonymous):

Yes so I got that part right. thank you. how do I write the sigma notation ?

OpenStudy (anonymous):

since you are close to zero, you can probably get a good estimation by taking \[\int_0^{.2}x^2dx\]

OpenStudy (anonymous):

the sixth power term will not contribute that much

OpenStudy (anonymous):

thanks